# acmart.bib

### v1

@@ 3.2. Training and Evaluation Configuration
For both tasks, we trained the circuits for 200 epochs with a batch size of 256, using the Adam optimizer with a learning rate of 0.01. The training objective was to minimize the loss between circuit predictions and the corresponding labels or target values in the training set. All training procedures were executed on noiseless simulators using the TorchQuantum (Wang_2022) framework, with computations performed on a 16-core AMD Ryzen Threadripper PRO 5955WX processor.
@@
(4) $\text{Score}(C)=\text{CNR}(C)^{\alpha_{\text{CNR}}}\times\text{RepCap}(C),$
@@
4. Results
Figure 2. Unbalance classification results. (a): Spearman correlation between original final score and test loss of Elivagar (10.1145/3620665.3640354). (b): Spearman correlation between revised final score and test loss.
Table 1. Average and standard deviation of selected metrics for PAMPA test set. The baseline for PRAUC is 0.808.
@@ 4. Results
MSE $0.7955\pm 0.1195$ $0.6002$ $0.1293\pm 0.0134$
Accuracy $0.7563\pm 0.0878$ $0.8250$ $0.8452\pm 0.0178$
F1 $0.8507\pm 0.0928$ $0.8966$ $0.9116\pm 0.0108$
PRAUC $0.8362\pm 0.0368$ $0.8801$ $0.9171\pm 0.0170$
@@
MSE $0.1201\pm 0.0112$ $0.1111\pm 0.0109$
Accuracy $0.8381\pm 0.0180$ $0.8569\pm 0.0164$
F1 $0.9101\pm 0.0108$ $0.9276\pm 0.0094$
PRAUC $0.9218\pm 0.0167$ $0.9195\pm 0.0178$
We evaluated our system’s performance using two primary metrics: the Spearman correlation between the final score and test MSE loss, and task-specific performance measures. For the classification task, we analyzed the Spearman correlation between the final score and PR-AUC, while for the regression task, we examined the Spearman correlation between the final score and the Spearman’s rank correlation coefficient of predictions versus true values. These ranking correlations are particularly crucial in ADME regression tasks.
4.1. Performance on Imbalanced Classification
Figure 2 compares the performance of the original Elivagar scoring system with our enhanced scoring method on the imbalanced classification task. As shown in Figure 2(a), the original Elivagar (10.1145/3620665.3640354) final score demonstrates a weak Spearman correlation (approximately -0.1) with the average test loss, indicating no significant predictive relationship between the scoring metric and actual circuit performance in imbalanced scenarios. In contrast, our revised scoring method achieves a moderate positive Spearman correlation of approximately 0.4 with the test loss, suggesting that our modifications successfully capture circuit performance characteristics in imbalanced classification tasks.
We also evaluate the performance of our quantum-generated circuits against classical machine learning models (XGBoost, Random Forest, SVM) on the PAMPA test set. As shown in Table 1, the QML circuits exhibit a consistent performance gap compared to classical baselines. For example, the best-performing QML circuit achieves an F1-score of 0.8966 and a PRAUC of 0.8801, while classical models such as XGBoost and SVM achieve superior results (F1: 0.9116–0.9276, PRAUC: 0.9171–0.9218). Notably, the average QML performance (e.g., MSE: 0.7955 $\pm$ 0.1195 vs. XGBoost: 0.1293 $\pm$ 0.0134) further underscores this gap. These results highlight the need for improved circuit generation strategies in our QCS framework, particularly to bridge the disparity in robustness and predictive power.
Figure 3. Regression results. (a): Spearman correlation between revised score and test loss of Elivagar (10.1145/3620665.3640354). (b): Spearman correlation between ranking relation of predicted and ground truth and revised score.
Table 2. Average and standard deviation of selected metrics for CaCo2 test set.
@@ 4.1. Performance on Imbalanced Classification
MSE $0.1376\pm 0.0336$ $0.0955$ $0.2421\pm 0.0242$
Spearman R $0.0929\pm 0.1386$ $0.4537$ $0.7433\pm 0.0359$
Metric Random Forest SVR
MSE $0.1977\pm 0.0177$ $0.3458\pm 0.0282$
Spearman R $0.7321\pm 0.0361$ $0.5606\pm 0.0564$
4.2. Regression Task Performance
Figure 3 illustrates our regression task results. Figure 3(a) shows the correlation between the revised final score and test loss, with a Spearman correlation coefficient of 0.18, indicating a weak positive correlation. Circuits with scores above 0.60 demonstrate consistently higher test losses. However, for circuits with scores below 0.6, the test loss clusters between 0.10 and 0.15, indicating limited discriminative power in this range.
We further analyzed the Spearman correlation between circuit predictions and ground truth values on the test set and the result is shown on figure 3(b) and table 2, as ranking accuracy is particularly relevant for medical applications. The near-zero correlation suggests that while our revised scoring method can identify poorly performing circuits, it has limited ability to predict how well a circuit will model the underlying distribution of the test set.
Finally, we evaluate the regression performance of our generated QML circuits against classical machine learning models (XGBoost, Random Forest, and SVR). As shown in Table 2, the QML circuits demonstrate superior performance, with the best-performing circuit achieving an MSE of 0.0955. This outperforms all classical baselines, including XGBoost (0.2421 $\pm$ 0.0242), Random Forest (0.1977 $\pm$ 0.0177), and SVR (0.3458 $\pm$ 0.0282). Even the average QML performance (0.1376 $\pm$ 0.0336) surpasses classical models, highlighting the robustness of our approach. These results underscore the potential of quantum-generated circuits in regression tasks and validate the effectiveness of the circuit generation strategy in our QCS framework.
5. Conclusion
Evaluation on medical task datasets demonstrates moderate correlation between our revised scoring method and final test loss. However, in regression tasks, while our approach successfully identifies low-performance circuits, it shows limitations in predicting circuits’ ability to correctly rank test set drugs—a crucial factor for medical applications. This limitation may stem from the disconnect between similarity-based scoring and distribution modeling capability.
Future work will focus on expanding validation across diverse ADME datasets and testing on physical quantum hardware. Additionally, we plan to develop alternative scoring mechanisms specifically tailored for regression tasks.
[…]

### v2

@@ 4.3 Training and Evaluation Configuration
For both tasks, we trained the circuits for 200 epochs with a batch size of 256, using the Adam optimizer with a learning rate of 0.01. The training objective was to minimize the loss between circuit predictions and the corresponding labels or target values in the training set. All training procedures were executed on noiseless simulators using the TorchQuantum [Wang_2022] framework, with computations performed on a 16-core AMD Ryzen Threadripper PRO 5955WX processor.
@@
$\text{Score}(C)=\text{CNR}(C)^{\alpha_{\text{CNR}}}\times\text{RepCap}(C),$ (7)
@@
Beside evaluating on classical computer, we also evaluate four classification tasks and two regression tasks’ circuits on real IBM Rensselaer quantum hardware. Both of these experiments results are shown in Section 5 and Appendix A.
5 Results
We evaluated performance using metrics tailored to each task. For both classification and regression, we calculated the Spearman correlation between the final score and the test Mean Squared Error. For classification tasks, we used Accuracy, F1-score, and the Area Under the Precision-Recall Curve (PR-AUC). For regression tasks, we additionally reported the MSE loss itself. All our implementation code can be found at https://github.com/zkysfls/quantum_admet.
5.1 Classification Tasks Performance
(a) QCS_ADME on BBB_Martins
(b) Elivagar on BBB_Martins
(c) QCS_ADME on Hia_Hou
(d) Elivagar on Hia_Hou
Figure 3: Spearman correlation between scores and test loss on unbalanced classification tasks BBB_Martins and Hia_Hou. Our revised score shows a stronger correlation compared to the original score.
Figure 4: The overview comparison between our proposed QCS method with other Quantum Circuit Search methods.
Figure 3 compares the performance of the original Elivagar scoring system with our enhanced scoring method on two imbalanced classification tasks: BBB_Martins and Hia_Hou. As shown in Figure 3(b) and Figure 3(d), the original Elivagar [10.1145/3620665.3640354] final score demonstrates a weak predictive relationship with the average test loss. On the BBB_Martins dataset, the correlation is negligible and not statistically significant ($R=0.0269,p=0.7520$), while on the Hia_Hou dataset, it is weak but significant ($R=0.2032,p=0.0015$). In contrast, our revised scoring method shows a consistently stronger correlation. On BBB_Martins (Figure 3(a)), it achieves a Spearman correlation of ($R=0.2244,p=0.0006$), and on Hia_Hou (Figure 3(c)), it reaches a moderate correlation of ($R=0.4121,p<0.0001$). These results suggest that our modifications successfully capture circuit performance characteristics in these two imbalanced classification tasks, making our score a more reliable metric. While Figure 3 illustrates two representative datasets where QCS-ADME strengthens the score–test-loss correlation, the effect is dataset-dependent. We therefore report correlations across all classification datasets in Appendix Table 18 and analyze failure cases where correlation improvements are weak or negative in section 5.4.
We also evaluate the performance of our quantum-generated circuits against other Quantum Circuit Search methods on all selected classification tasks. For this comparison, we choose the QML circuits ranked in the Top-25 by our revised score. As shown in Figure 4, we compare four circuit selection methods: our proposed QCS (orange), Elivagar [10.1145/3620665.3640354] (blue), QuantumNAS [Wang_2022] (red). Across all three classification metrics, a clear trend emerges, with our proposed QCS method consistently demonstrating robust and high-quality performance. For a majority of the tasks—including CYP2C9_Substrate_CarbonMangels, CYP2D6_Substrate_CarbonMangels, BBB_Martins, hERG, and HIA_Hou—the circuits selected by QCS achieve results that are either the best or highly competitive with the other leading methods. For instance, on the BBB_Martins task, QCS leads across all three metrics, achieving an accuracy of approximately 86%, an F1-score of approximately 92%, and a PR-AUC of approximately 92%.
Furthermore, we evaluated the performance of our quantum-generated circuits against classical machine learning models (XGBoost, Random Forest, SVM, and Logistic Regression). As shown in Table 1 and Table 2, the performance of the QML circuits presents a mixed picture when compared to the classical baselines. For the BBB_Martins task, the best-performing QML circuit achieves a superior F1-score of 0.9198, outperforming all listed classical models. However, its PR-AUC of 0.9301 lags behind the classical approaches, such as Random Forest (0.9650) and XGBoost (0.9639). On the Hia_Hou task, the performance gap is more pronounced. The best QML circuit’s F1-score (0.8913) is surpassed by several classical models like XGBoost (0.9554), and its PR-AUC (0.9384) is substantially lower than the near-perfect scores achieved by all the classical baselines (all >0.995>0.995). These results, with detailed comparisons for additional tasks available in the Appendix, highlight that while QML circuits can be competitive on certain metrics, a significant gap often remains, particularly in PR-AUC.
Table 1: Performance comparison for BBB_Martins.
@@ 5.1 Classification Tasks Performance
Accuracy $0.7831\pm 0.0643$ $0.8596$ $0.8794\pm 0.0158$
F1 $0.8684\pm 0.0503$ $0.9198$ $0.8798\pm 0.0158$
PR-AUC $0.8804\pm 0.0221$ $0.9301$ $0.9639\pm 0.0104$
Metric Random Forest SVM Logistic Regression
Accuracy $0.8871\pm 0.0157$ $0.8780\pm 0.0160$ $0.8578\pm 0.0179$
F1 $0.8795\pm 0.0176$ $0.8659\pm 0.0187$ $0.8599\pm 0.0176$
PR-AUC $0.9650\pm 0.0114$ $0.9578\pm 0.0118$ $0.9464\pm 0.0154$
Table 2: Performance comparison for Hia_Hou.
Metric QML average QML best XGBoost
Accuracy $0.6971\pm 0.0830$ $0.8291$ $0.9573\pm 0.0182$
F1 $0.7980\pm 0.0733$ $0.8913$ $0.9554\pm 0.0196$
PR-AUC $0.8529\pm 0.0552$ $0.9384$ $0.9976\pm 0.0018$
Metric Random Forest SVM Logistic Regression
Accuracy $0.9056\pm 0.0261$ $0.8816\pm 0.0298$ $0.9408\pm 0.0218$
F1 $0.8954\pm 0.0310$ $0.8644\pm 0.0367$ $0.9393\pm 0.0228$
PR-AUC $0.9954\pm 0.0031$ $0.9964\pm 0.0022$ $0.9960\pm 0.0027$
5.2 Regression Task Performance
Figure 5 showcases the relationship between our proposed circuit scores and final model performance, measured by average test loss, across four distinct regression tasks. The scatter plots for the VDss_Lombardo, Half_Life_Obach, Clearance_Hepatocyte_AZ, Clearance_Microsome_AZ tasks consistently demonstrate a statistically significant negative correlation between the two variables. This is the desired outcome, as it confirms that circuits with higher scores are strongly associated with lower test losses, indicating better predictive performance. The Spearman correlation coefficients (R) are moderate and negative, ranging from -0.3367 to -0.4881, and are all highly significant ($p<0.0001$). These results validate our scoring metric as an effective tool for identifying superior quantum circuits for regression problems.
(a) Half_Life_Obach
(b) Clearance_Hepatocyte_AZ
(c) Clearance_Microsome_AZ
(d) VDss_Lombardo
Figure 5: Spearman correlation between scores and test loss on regression tasks.
Figure 6: The comparison between our proposed QCS method with classical ML models on regression tasks
Finally, we evaluate the regression performance of our generated QML circuits against classical machine learning models. We follow the classification comparsion setting that take the Top-25 circuits based on our score and compare to the same set of classical ML models (XGBoost, Random Forest, and SVM). The results in figure 6 reveal a significant performance gap, with the classical models consistently achieving substantially lower MSE values than the QML approach in all evaluated cases. For instance, on the Clearance_Hepatocyte_AZ and Clearance_Microsome_AZ tasks, the QML model’s MSE of 0.455 and 0.420, respectively, is markedly higher than its classical counterparts. This disparity becomes even more pronounced on the Half_Life_Obach and VDss_Lombardo tasks, where classical models achieve exceptionally low MSEs ( 0.011 for Random Forest and 0.003 for SVM), while the QML model’s loss remains significantly higher at 0.279 and 0.307, respectively. These findings underscore a key challenge: while our framework can effectively generate and evaluate quantum circuits, bridging the performance gap to established classical regression algorithms remains a critical area for future research.
5.3 Real hardware Performance
Figure 7: The comparison between simulation and real hardware performances of selected classification tasks
To further investigate our proposed method’s performance, we chose four classification tasks and two regression tasks to run on IBM Rensselaer, which is the hardware configuration we chose to generate our circuits. Figure 7 shows the overview comparison of simulation to real hardware on classification tasks. We could see that a general degradation in performance is observed when transitioning from simulation to noisy hardware execution. This is particularly evident in the BBB_Martins task, where the average accuracy drops from 0.783 to 0.494 and the F1-score decreases from 0.868 to 0.538. This trend, coupled with a significant increase in performance variance on the hardware, highlights the impact of device noise. However, the Bioavailability_Ma task presents a compelling and counter-intuitive exception. On this task, the circuits executed on real hardware consistently outperformed their simulated counterparts across all metrics. The average PR-AUC, for instance, improved from 0.523 in simulation to 0.770 on the hardware, while the Mean Squared Error advantageously decreased from 1.042 to 0.931. This suggests that for certain problem landscapes, the inherent noise of the hardware may act as a form of regularization, potentially aiding the optimization process and leading to superior generalization compared to the ideal, noiseless environment.
Figure 8: The comparison between simulation and real hardware performances of selected regression tasks
For regression tasks, figure 8 shows an overview comparison. The result reveals a significant performance gap between noiseless simulation and execution on real hardware. This ”sim-to-real” gap is evident in both the average and best-case scenarios. The mean loss increased substantially by 24.9% for the Hepatocyte dataset and 37.5% for the Microsome dataset when moving from simulation to hardware. Similarly, the best-case performance was also degraded, with the minimum achievable loss increasing by 4.8% and 11.4% for the respective datasets. Beyond the degradation in accuracy, the performance on real hardware demonstrated markedly higher instability. This is quantified by the standard deviation of the loss, which was over five times larger on hardware for both datasets, and is visually confirmed by the wider performance distributions in the box plots. These results underscore the profound impact of hardware noise, which not only diminishes the predictive power of the models but also severely impacts the reliability and consistency of their performance on regression tasks.
In summary, our ”sim-to-real” analysis demonstrates that hardware noise has a complex and task-dependent impact on performance. While noise was uniformly detrimental to our regression tasks, consistently increasing error and instability, its effect on classification was more nuanced. We found that for some classification tasks, noise paradoxically improved results, suggesting it can act as a potential regularizer in addition to being a performance bottleneck. This dual nature of hardware noise necessitates task-aware strategies for developing effective near-term quantum algorithms.
5.4 Ablation Study and Failure Analysis
To justify the design of the weighted matrix $R_{w}$, we conducted an ablation study comparing our method against the baseline Elivagar scoring which lacks imbalance adaptation. As shown in Table 18, the impact of the weighting mechanism is task-dependent.
Success Cases: On the hERG and Bioavailability_Ma datasets, the introduction of $R_{w}$ transformed the scoring metric from random or weakly predictive to strongly predictive. For instance, on hERG, the Spearman correlation improved from $R=0.058$ to R=0.276R=-0.276, and the best-found circuit achieved an F1-score of 0.852, surpassing the baseline’s 0.843. This demonstrates that for datasets where the minority class features are distinct but underrepresented, the weighted RepCap successfully guides the search toward robust architectures.
textbfLimitations and Failure Modes: Conversely, we acknowledge specific limitations regarding data quality and study scope. First, we observed inconsistencies in datasets such as HIA_Hou and BBB_Martins, where the metric exhibited a positive correlation with test loss (e.g., $R=0.412$ for HIA_Hou). We hypothesize this is due to the noise profile of these specific ADME datasets. In scenarios where the minority class contains significant label noise or outliers, the weighted metric effectively rewards overfitting to the training geometry, which in result prioritize circuits that separate these outliers at the expense of learning the true decision boundary. This suggests that while $R_{w}$ is a powerful tool for imbalance, it requires datasets with reliable minority labels to function optimally. Second, we restricted our evaluation to TDC benchmarks rather than larger industrial datasets due to the computational constraints of simulating Quantum Circuit Search. The processing of industrial-scale feature spaces currently exceeds NISQ-era simulation capacities, necessitating future hardware advancements to bridge this gap.
6 Conclusion
Evaluation on medical task datasets demonstrates moderate correlation between our revised scoring method and final test loss. This validates our score as a more reliable predictor of model performance than previous methods, particularly for regression tasks where it showed consistent negative correlation. Circuits selected by our framework were highly competitive with those from other state-of-the-art quantum circuit search methods. However, when compared to classical models, a performance gap remains; while our QML circuits were competitive in some classification metrics, they were significantly outperformed in regression tasks. Finally, our ”sim-to-real” analysis revealed that the impact of hardware noise is highly task-dependent: it was uniformly detrimental to regression performance but had a more nuanced effect on classification, even paradoxically improving results in one instance. These findings highlight both the utility of our adapted framework and the ongoing challenges of bridging the performance gap to classical algorithms and navigating the complex effects of noise in the NISQ era.
There are several possible directions to further improve our work. Firstly, we will focus on developing advanced methods to generate more competitive candidate circuits. Second, while our use of Clifford Noise Resilience (CNR) effectively guides the search toward robust architectures, the sim-to-real gap in regression tasks suggests the need for further mitigation. Future development could incorporate task-aware circuit pruning, where gates with negligible gradients regarding the specific regression loss are iteratively removed before hardware execution. Additionally, we plan to incorporate label-noise robustness terms into the $R_{w}$ formulation.
7 Declarations
Funding: This work is funded through the IBM-RPI Future of Computing Research Collaboration.
Clinical trial number: not applicable.
Appendix A Additional results
Table 3: Performance comparison for CYP2C9_Substrate_CarbonMangels task.
Metric QML average QML best XGBoost
Accuracy $0.6684\pm 0.0590$ $0.7407$ $0.7041\pm 0.0391$
F1 $0.2630\pm 0.1373$ $0.4691$ $0.6590\pm 0.0472$
PR-AUC $0.3390\pm 0.0426$ $0.4297$ $0.3922\pm 0.0739$
Metric Random Forest SVM Logistic Regression
Accuracy $0.6971\pm 0.0392$ $0.7193\pm 0.0374$ $0.7110\pm 0.0385$
F1 $0.6033\pm 0.0501$ $0.6024\pm 0.0494$ $0.6633\pm 0.0473$
PR-AUC $0.3760\pm 0.0670$ $0.3921\pm 0.0682$ $0.3854\pm 0.0723$
Table 4: Performance comparison for CYP2D6_Substrate_CarbonMangels task.
Metric QML average QML best XGBoost
Accuracy $0.6871\pm 0.0272$ $0.7333$ $0.7276\pm 0.0379$
F1 $0.2574\pm 0.1396$ $0.5405$ $0.7229\pm 0.0391$
PR-AUC $0.4189\pm 0.0489$ $0.5174$ $0.5757\pm 0.0768$
Metric Random Forest SVM Logistic Regression
Accuracy $0.7407\pm 0.0383$ $0.7405\pm 0.0384$ $0.7190\pm 0.0396$
F1 $0.7227\pm 0.0427$ $0.7191\pm 0.0437$ $0.7151\pm 0.0410$
PR-AUC $0.6292\pm 0.0744$ $0.6390\pm 0.0730$ $0.5690\pm 0.0803$
Table 5: Performance comparison for DILI task.
Metric QML average QML best XGBoost
Accuracy $0.6279\pm 0.0534$ $0.7292$ $0.7992\pm 0.0401$
F1 $0.6221\pm 0.0725$ $0.7327$ $0.7966\pm 0.0411$
PR-AUC $0.6811\pm 0.0757$ $0.8111$ $0.8783\pm 0.0374$
Metric Random Forest SVM Logistic Regression
Accuracy $0.8218\pm 0.0380$ $0.8115\pm 0.0384$ $0.8019\pm 0.0399$
F1 $0.8195\pm 0.0388$ $0.8094\pm 0.0393$ $0.8009\pm 0.0403$
PR-AUC $0.8988\pm 0.0365$ $0.8992\pm 0.0332$ $0.8479\pm 0.0465$
Table 6: Performance comparison for hERG task.
Metric QML average QML best XGBoost
Accuracy $0.7206\pm 0.0330$ $0.7576$ $0.8177\pm 0.0323$
F1 $0.8244\pm 0.0288$ $0.8546$ $0.8039\pm 0.0367$
PR-AUC $0.7953\pm 0.0427$ $0.8584$ $0.9175\pm 0.0280$
Metric Random Forest SVM Logistic Regression
Accuracy $0.8330\pm 0.0325$ $0.7947\pm 0.0350$ $0.7794\pm 0.0347$
F1 $0.8142\pm 0.0386$ $0.7695\pm 0.0421$ $0.7478\pm 0.0427$
PR-AUC $0.9130\pm 0.0281$ $0.8968\pm 0.0342$ $0.8343\pm 0.0443$
Table 7: Performance comparison for Pgp_Broccatelli task.
Metric QML average QML best XGBoost
Accuracy $0.6150\pm 0.0502$ $0.7061$ $0.8207\pm 0.0249$
F1 $0.6396\pm 0.0554$ $0.7273$ $0.8207\pm 0.0249$
PR-AUC $0.6596\pm 0.0750$ $0.7704$ $0.9066\pm 0.0200$
Metric Random Forest SVM Logistic Regression
Accuracy $0.8292\pm 0.0235$ $0.8292\pm 0.0237$ $0.8044\pm 0.0256$
F1 $0.8291\pm 0.0235$ $0.8291\pm 0.0237$ $0.8044\pm 0.0256$
PR-AUC $0.9173\pm 0.0181$ $0.9142\pm 0.0183$ $0.8889\pm 0.0216$
Table 8: Performance comparison for Clearance_Hepatocyte_AZ task.
Metric QML average QML best XGBoost
MSE $0.4552\pm 0.0159$ $0.4190$ $0.5335\pm 0.0491$
Spearman -- 0.3307-0.3307 $0.2184\pm 0.0589$
@@ Appendix A Additional results
MSE $0.4467\pm 0.0405$ $0.4123\pm 0.0461$
Spearman $0.2479\pm 0.0597$ $0.3327\pm 0.0544$
Table 9: Performance comparison for Clearance_Microsome_AZ task.
@@
MSE $0.4202\pm 0.0297$ $0.3470$ $0.2775\pm 0.0401$
Spearman -- 0.4406-0.4406 $0.4643\pm 0.0565$
Metric Random Forest SVM
MSE $0.2756\pm 0.0387$ $0.2848\pm 0.0433$
Spearman $0.4870\pm 0.0524$ $0.4571\pm 0.0535$
Table 10: Performance comparison for Half_Life_Obach task.
Metric QML average QML best XGBoost
MSE $0.2788\pm 0.0757$ $0.1262$ $0.0197\pm 0.0132$
Spearman -- 0.4703-0.4703 $0.1736\pm 0.0945$
Metric Random Forest SVM
MSE $0105\pm 0.0052 0.0$$0063\pm 0.0006$
Spearman $0.3450\pm 0.0866$ $0.1401\pm 0.0870$
Table 11: Performance comparison for VDss_Lombardo task.
Metric QML average QML best XGBoost
MSE $0.3073\pm 0.0521$ $0.1412$ $0.0141\pm 0.0128$
Spearman -- 0.4651-0.4651 $0.2883\pm 0.0674$
Metric Random Forest SVM
MSE $0\pm 0.0024 0.0033±0.00$$3\pm 0.0002$
Spearman $0.3992\pm 0.0600$ $0.2056\pm 0.0649$
Table 12: Real Hardware vs Simulation comparison for CYP2C9_Substrate_CarbonMangels.
Metric Simulation Max Simulation Mean Real HW Max
Accuracy $0.741$ $0.668\pm 0.059$ $0.719$
F1 $0.469$ $0.263\pm 0.137$ $0.455$
PR-AUC $0.430$ $0.339\pm 0.043$ $0.343$
MSE $1.013$ $0.929\pm 0.049$ $1.038$
Metric Real HW Mean
Accuracy $0.541\pm 0.098$
F1 $0.293\pm 0.111$
PR-AUC $0.278\pm 0.035$
MSE $0.988\pm 0.037$
Table 13: Real Hardware vs Simulation comparison for CYP2D6_Substrate_CarbonMangels.
Metric Simulation Max Simulation Mean Real HW Max
Accuracy $0.733$ $0.687\pm 0.027$ $0.704$
F1 $0.540$ $0.257\pm 0.140$ $0.497$
PR-AUC $0.517$ $0.419\pm 0.049$ $0.444$
MSE $0.972$ $0.887\pm 0.041$ $1.159$
Metric Real HW Mean
Accuracy $0.576\pm 0.097$
F1 $0.314\pm 0.134$
PR-AUC $0.347\pm 0.048$
MSE $0.979\pm 0.063$
Table 14: Real Hardware vs Simulation comparison for BBB_Martins.
Metric Simulation Max Simulation Mean Real HW Max
Accuracy $0.860$ $0.783\pm 0.064$ $0.800$
F1 $0.920$ $0.868\pm 0.050$ $0.889$
PR-AUC $0.930$ $0.880\pm 0.022$ $0.891$
MSE $0.973$ $0.806\pm 0.105$ $1.234$
Metric Real HW Mean
Accuracy $0.494\pm 0.205$
F1 $0.538\pm 0.289$
PR-AUC $0.831\pm 0.027$
MSE $0.992\pm 0.132$
Table 15: Real Hardware vs Simulation comparison for Bioavailability_Ma.
Metric Simulation Max Simulation Mean Real HW Max
Accuracy $0.635$ $0.540\pm 0.034$ $0.758$
F1 $0.711$ $0.658\pm 0.043$ $0.862$
PR-AUC $0.668$ $0.523\pm 0.062$ $0.851$
MSE $1.126$ $1.042\pm 0.046$ $1.029$
Metric Real HW Mean
Accuracy $0.615\pm 0.111$
F1 $0.724\pm 0.131$
PR-AUC $0.770\pm 0.042$
MSE $0.931\pm 0.079$
Table 16: Real Hardware vs Simulation comparison for Clearance_Hepatocyte_AZ.
Metric Simulation Real Hardware Difference (%)
Mean Loss $0.4552\pm 0.0156$ $0.5685\pm 0.0874$ $24.9$
Min Loss $0.4190$ $0.4392$ $4.8$
Max Loss $0.4745$ $0.7222$
Median $0.4519$ $0.5490$
Table 17: Real Hardware vs Simulation comparison for Clearance_Microsome_AZ dataset.
Metric Simulation Real Hardware Difference (%)
Mean Loss $0.4202\pm 0.0291$ $0.5779\pm 0.1505$ $37.5$
Min Loss $0.3470$ $0.3867$ $11.4$
Max Loss $0.4570$ $0.8707$
Median $0.4265$ $0.4955$
Table 18: Correlation summary across classification datasets. Spearman $R$ is computed between circuit score and test loss (MSE). We also report the best F1 achieved among evaluated circuits for each scoring method.
Baseline (Elivagar) Ours (QCS-ADME)
Dataset Spearman $R$ F1 (best) Spearman $R$ F1 (best)
hERG 0.058 0.843 -0.276 0.852
BBB_Martins 0.021 0.916 0.224 0.920
Bioavailability_Ma -0.182 0.678 -0.427 0.711
HIA_Hou 0.203 0.887 0.412 0.891
CYP2C9_Substrate_CarbonMangels 0.218 0.286 0.340 0.286
CYP2D6_Substrate_CarbonMangels 0.031 0.471 -0.136 0.541
DILI -0.339 0.756 -0.159 0.705
Pgp_Broccatelli -0.162 0.752 -0.052 0.719
Statistically significant correlation ($p<0.05$).